OpenStudy (anonymous):

What is the best way to classify each equation? 1. 8x+24=2(4x+12) 2. 5x+18-x=2(2x+8) 3. 7(3x-2)=20x-13 4. 3x+2(x-10)=5(x-4) Give each one of the numbers a letter: A. Identify B. Contradiction C. Neither

OpenStudy (anonymous):

\[ 8x+24=2(4x+12)\]if both sides are equal, it is an identity lets check by the distributive law \[ 8x+24=2(4x+12)\\ 8x+24=8x+24\] yup they are equal

OpenStudy (anonymous):

Identity

OpenStudy (anonymous):

\[5x+18-x=2(2x+8)\] multiply out and get \[5x+18-x=4x+16\] combine like terms andget \[4x+18=4x+16\]

OpenStudy (anonymous):

seems unlikely that you could add 18 to a number, add 16 to that same number, and get the same answer that makes this a "contradiction"

OpenStudy (amistre64):

if f(x) = g(x) is true for all x, then we have an identity if f(x) = g(x) is false for all x, then we have a contradiction if f(x) = g(x) is true for only a few specific x, then we neither an id nor a contra

OpenStudy (anonymous):

@amistre64 By the way that dosent help because i dont know how to solve for X in the first place

OpenStudy (anonymous):

@iambatman help?